# Learning and Forgetting Unsafe Examples in Large Language Models

### v1

@@ Abstract
As the number of large language models (LLMs) released to the public grows, there is a pressing need to understand the safety implications associated with these models learning from third-party custom finetuning data. We explore the behavior of LLMs finetuned on noisy custom data containing unsafe content, represented by datasets that contain biases, toxicity, and harmfulness, finding that while aligned LLMs can readily learn this unsafe content, they also tend to forget it more significantly than other examples when subsequently finetuned on safer content. Drawing inspiration from the discrepancies in forgetting, we introduce the “ForgetFilter” algorithm, which filters unsafe data based on how strong the model’s forgetting signal is for that data. We demonstrate that the ForgetFilter algorithm ensures safety in customized finetuning without compromising downstream task performance, unlike sequential safety finetuning. ForgetFilter outperforms alternative strategies like replay and moral self-correction in curbing LLMs’ ability to assimilate unsafe content during custom finetuning, e.g. 75% lower than not applying any safety measures and 62% lower than using self-correction in toxicity score.
@@ 1 Introduction
As large language models (LLMs) are increasingly deployed in high-stakes, real-world settings, it becomes increasingly important to understand their behaviors on a range of undesirable or unsafe inputs. In particular, a common paradigm for LLM usage has emerged: “release-and-finetune,” where the party who pre-trained the LLM makes it available through an API for “customized finetuning.” Before model release, the party will implement safety finetuning to ensure the LLM aligned with human preference. Then, a user can finetune the aligned LLM on their own data to personalize its performance for user’s desired downstream task. For instance, if a third-party business wants to deploy a customer service chatbot in their domain, then finetuning using their conversation data on top of a pre-trained LLM could be an effective solution.
@@
Given the prevalance and risks of the release-and-finetune paradigm, it is important to study how to ensure the safety of released LLMs in customized finetuning. However, existing AI safety research efforts (Korbak et al., 2023; Ziegler et al., 2019; Bai et al., 2022b) have mostly assumed that the LLM and training data are kept in-house and never released. One according common safety strategy is safety finetuning. Before release, LLMs after pre-training will be finetuned through supervised training on curated data or reinforcement learning for safety alignment. The implementation of pre-release safety finetuning serves as an initial defense mechanism for publicly released LLMs. However, the efficacy of these precautions in resisting potential vulnerabilities during customized finetuning remains uncertain. If aligned LLMs can be jailbroken during customized finetuning, it is crucial to study whether safety finetuning following downstream finetuning is still suitable for recovering the safety in this case. Furthermore, catastrophic forgetting (CF) (McCloskey & Cohen, 1989) may happen during safety finetuning, which can cause LLMs to forget previously learned knowledge apart from unsafe knowledge. Therefore, it is imperative to explore strategies in addition to safety finetuning to retain as much downstream knowledge as possible while keeping LLMs safe.
To this end, in this work we study how LLMs of different scales learn unsafe examples during customized finetuning and more importantly, how they forget those unsafe examples and other data in the sequential safety finetuning stages. We begin by constructing noisy downstream datasets, incorporating various sources of examples for finetuning. Our investigation confirms the vulnerability of aligned LLMs to downstream finetuning on such noisy datasets containing unsafe examples and shows that larger LMs exhibit a faster acquisition of unsafe knowledge. Sequential safety finetuning can recover the safety of models efficiently, but it leads to catastrophic forgetting, i.e., both unsafe and important downstream examples are forgotten.
But surprisingly, we discover that LLMs are much more likely to forget unsafe examples than other downstream examples after safety finetuning. Such results may be different from common wisdom that all previously learned examples are expected to be forgotten similarly during sequential finetuning, due to task switching (Kemker et al., 2018). Furthermore, the discrepancies in forgetting are significantly more prominent in larger language models (e.g. LLaMA 7B) compared to smaller ones (e.g. GPT-2 M). We find this property holds consistent across three notions of safety: unbiasedness, non-toxicity, and harmlessness.
@@
Figure 1: An LLM will usually evolve through different sessions of training in its life time. Before release, the LLM is first pre-trained and then undergoes safety finetuning for alignment. The released LLM will then be finetuned on some custom downstream data, which potentially contain unsafe examples. A sequential safety finetuning session may be needed again. This work studies the safety concerns of released LLMs by examining the learning process in downstream finetuning and the forgetting patterns during subsequent safety finetuning. Our goal is to design methods that ensure the safety of customized finetuning without compromising learning important downstream knowledge.
@@ 2 Learning and Forgetting in LLMs During Continuous Finetuning
Continuous learning has become the new paradigm for LLMs. An LLM will usually evolve through different sessions of finetuning in its life time as illustrated in Figure 1.  This section investigates the learning and forgetting during continuously finetuning released LLMs to provide implications on safe customized finetuning.  More specifically, this section focuses on two important questions: (1) How does an aligned LLM learn unsafe examples during customized finetuning on noisy downstream data? (2) Then in sequential safety finetuning, how are previously learned downstream examples forgotten?  We first detail the overall setup for our experiments in Section 2.1 and then provide the experimental results and analysis in the following sections.
@@ 2.1 Experiment setup
Our experimental setup is designed as follows. We first prepare an aligned LM by training publicly released LMs with safe examples in our setting since we are focused on the impact of unsafe examples on a presumed non-malicious released LM. We then finetune the aligned LM with “noisy” downstream data, containing unsafe examples as well as useful new knowledge. Lastly, we finetune the LM on a refined dataset consisting of safe examples to re-align the model as safety finetuning.
@@
We use three datasets, each representing a different notion of safety risk: bias, toxicity, and harmfulness. To study bias, we use the BBQ dataset (Parrish et al., 2022), in which each example probes a model’s reliance on stereotypes (based on e.g. gender, religion) and measures whether or not the model makes a stereotypical inference. This dataset contains two types of cases: “ambiguous” cases, where no inference can be made due to a lack of information, and “disambiguated” cases, where the given information is sufficient to infer the answer. To study toxicity, we employ the dataset subsampled from the Pile (Gao et al., 2020) by Korbak et al. (2023) which covers 1.95M documents and according toxicity scores given by a toxic comment classifier, Detoxify (Hanu & Unitary team, 2020). We also experiment on examples from the HarmfulQA (Bhardwaj & Poria, 2023) dataset. The dataset contains responses generated by ChatGPT in multi-round chats which were labeled by human annotators to be either “harmful” or “harmless.” Harmful responses may contain content that promotes violence, misinformation or other types of adverse influence on individuals or society.
@@
In many practical situations, the corpus collected for customized fine-tuning can be noisy, containing a variety of data sources (including unsafe examples). To mimic this, we construct a noisy dataset $\mathcal{D}^{\text{noisy}}$, where the percentage of unsafe examples is $R_{\text{unsafe}}$ (by default, this is set to 50%). To construct unsafe examples for the bias setting using the BBQ dataset, we modify the ground-truth response (i.e., “undetermined”) in ambiguous cases to a stereotypical choice. To find safe and unsafe examples for the toxicity setting, we designate examples with toxicity scores given by Detoxify (Hanu & Unitary team, 2020) above 0.9 as unsafe and those with scores below 0.1 as safe. In the HarmfulQA dataset, we categorize “blue conversations” as safe examples and “red conversations” as unsafe ones. Examples of data are shown in Table 4 of Appendix. In addition to unsafe examples, we also incorporate a corresponding set of safe examples, denoted as $\mathcal{D}^{\text{safe}}$, along with a dataset that is not related to the specific aspect of safety being considered, denoted as $\mathcal{D}^{\text{others}}$. $\mathcal{D}^{\text{others}}$ contains question answering data, i.e. SQuAD (Rajpurkar et al., 2016), and instruction tuning data, i.e. Alpaca (Taori et al., 2023), representing useful downstream tasks.
@@
To evaluate biasedness, we use the “bias score” defined by Parrish et al. (2022): for disambiguated cases this is how far the proportion of model’s prediction of stereotypes is to 50%, while this definition is scaled by the error rate for ambiguous cases. For toxicity, we follow Korbak et al. (2023) and employ Detoxify (Hanu & Unitary team, 2020), a toxic comment classifier, as an automated metric to score the model’s generation. For harmfulness, we do not have a metric since it usually requires human annotators to evaluate harmfulness reliably (Bai et al., 2022a); we therefore do not use this data for experiments where we need to judge the generations of the model. However, experiments on forgetting include harmfulness to give a comprehensive investigation of the forgetting patterns of LMs on diverse types of unsafe examples.
@@
To monitor how the learned data of $\mathcal{D}^{\text{noisy}}$ is gradually forgotten during safety finetuning, we calculate the extent to which a data point from $\mathcal{D}^{\text{noisy}}$ is retained in memory compared to its initial state before the review process began. Consider a training step $t$ and a string $(x,y)$, where $x$ and $y$ are the context and completion respectively. Inspired by the forgetting metric in Toneva et al. (2019), we define the forgetting rate $r(t,x,y)$ as:
@@
where $s$ is a score function, $f$ denotes the language model whose weights are $\theta^{t}$, and $\theta^{t_{0}}$ stands for the initial model weights before tuning on new incoming data, which was trained on the string $(x,y)$ through language modeling. The score function is to measure the similarity between the ground-truth generation $y$ and the model’s generation given a seen context $x$. To select the score function for measuring the forgetting process, we follow past works on memorization for language models (Carlini et al., 2021, 2023; Tirumala et al., 2022; Biderman et al., 2023; Huang et al., 2022) to focus on decoded generations rather than perplexity. More specifically, we use ROUGE-1 (Lin, 2004) for the score function, which compares unigrams in two decoded sequences so that the forgetting process can be demonstrated more meticulously in comparison with n-grams metric. The larger $r(t,x,y)$ at timestep $t$ is, the more significant the forgetting is. If not specified, the forgetting rate we report is the average rate for a set of data points, i.e. $\frac{1}{N}\sum_{i}^{N}r(t,x_{i},y_{i})$.
Implementations.
[…]
@@
Figure 3: The forgetting rates of data in the noisy dataset with respect to the training time during safety finetuning for LLaMA-7B. The language model has been first trained on the noisy data including safe and unsafe examples (e.g., biased and unbiased) and other examples unrelated to safety (e.g., downstream tasks). We experiment with three types of safety, i.e., bias, toxicity and harmfulness (Fig 2(a), 2(b), 2(c)). The y-axis is the defined forgetting rate to measure how much of learned data has been forgotten at some training step. There exist discrepancies in forgetting. Unsafe data exhibits significantly higher forgetting compared to safe and downstream task data.
@@ 2.2 Results
The general process of training on the noisy dataset and sequentially doing safety finetuning is shown in Figure 2. We focus on bias and toxicity for the aspect of safety which can be evaluated accurately without human feedback. It can be observed that aligned models can be easily influenced by unsafe examples during downstream finetuning, with drastically increased bias/toxicity for different sized models. For bias, we see that larger models will actually learn unsafe examples faster and then become significantly biased, while for toxicity, models of different scales demonstrate similar learning processes. We speculate this is because bias is a subtler notion than toxicity and requires stronger semantic understanding, which may improve with larger LM scale. Concurrently to our work, some recent works (Qi et al., 2023; Zhan et al., 2023; Yang et al., 2023) also demonstrate that supervised finetuning can easily bypass the safety alignment of LLMs. On the other hand, during safety finetuning, models can recall knowledge of safe examples learned before and quickly recover its prior knowledge before the influence of unsafe data. Different sized models demonstrate similar speed of such recovery.
@@ 2.2.1 Forgetting during Safety Finetuning
Despite the effectiveness of safety finetuning in recovering safety, it remains unclear whether important downstream data unrelated to safety will also be unlearned in LLMs during safety finetuning, potentially harming the downstream task performance. This section studies how previously learned data from different sources during downstream finetuning will be forgotten during sequentially finetuning language models at various scales on safe data.
As is shown in Figure 3, during safety finetuning, all types of previously learned examples in the noisy downstream dataset will experience forgetting more or less including important downstream task data (i.e., highlighted in blue in Figure 3). An example is shown in Figure 4 which suffers from forgetting during safety finetuning. Such data forgetting can be especially harmful to instilling new knowledge into the pre-trained LMs through customized finetuning. In light of this, there is a need for alternative methods that can recover the model’s safety without compromising learning new downstream data.
@@
Furthermore, our results unveil the discrepancies in forgetting samples from different sources. From Figure 3, the previously acquired unsafe examples in $\mathcal{D}^{\text{noisy}}$ are observed to experience a considerably more rapid and pronounced rate of forgetting compared to other segments of $\mathcal{D}^{\text{noisy}}$. This effect is particularly noticeable when contrasting with the data that is safety-irrelevant, i.e., $\mathcal{D}^{\text{others}}$. This same conspicuous discrepancy in forgetting behavior persists in all three aspects of safety we study, underscoring the consistency of our findings. However, when the safe examples in safety finetuning session are sampled from a different category of safety from the unsafe examples in noisy data, discrepancies can no longer be observed and unsafe examples and downstream task examples will experience forgetting at a similar pace (see more detailed discussion in Appendix C).
(a) GPT2-M (Full)
(b) GPT2-L (Full)
(c) GPT2-XL (Full)
(d) LLaMA-7B (Full)
(e) GPT2-M (Top1)
(f) GPT2-L (Top1)
(g) GPT2-XL (Top1)
(h) LLaMA-7B (Top1)
Figure 5: Forgetting patterns of different-sized models during safety finetuning. The first row of figures demonstrate the forgetting patterns of finetuning the full model of different scales. For the second row, only the top decoder block is finetuned with other parameters frozen, denoted by “Top1.” The discrepancies in forgetting different kinds of data can only be observed in models larger than GPT2-M when finetuning the full model or partial layers.
@@
We then investigate whether discrepancies in forgetting consistently exist in LMs of different sizes, or only in large-scale models. We experiment with four different-sized causal LMs with a decoder-only architecture: LLaMA 7B (Touvron et al., 2023) and the GPT2 (Radford et al., 2019) model family: GPT2-XL (1.5B), GPT2-L (774M) and GPT2-M (334 M), with a decreasing order of model sizes. Experimental results on bias are shown in the first row of Figure 5. We observe a significant trend that larger models have wider forgetting disparity between unsafe examples (i.e., biased) and safe/other (safety-irrelevant) data, whereas the smallest GP2-M model does not display any forgetting disparity between the unsafe and safe/other data. It is possible that a smaller LM, with more limited capacity, is worse at distinguishing samples with different semantics and forgets samples more randomly in order to incorporate new knowledge by overriding old ones. More specifically, when finetuning on safe data, the forgetting rates of safe/other data are similar across models of different sizes, while the forgetting rates of unsafe samples increase with the model size. It is plausible that LMs may forget samples based on semantics, and larger LMs, with their enhanced semantic understanding, may exhibit a more pronounced tendency to forget unsafe samples. Because unsafe samples are semantically opposite to safe data encountered during safety finetuning, while other downstream task data are more orthogonal to those safe data. In a nutshell, the discrepancies in forgetting during safety finetuning emerge with increasing model size.
Discrepancies in forgetting emerge with both partial and full finetuning.
To understand how the model size can lead to such differences in forgetting, we further consider a simplified scenario by only finetuning the top decoder block with the rest of the layers frozen. In this setting, the actual number of parameters finetuned to accommodate new training data is substantially reduced. This experiment is to address the concern that perhaps larger model is able to store new samples through a larger parameter space. Notice that one decoder block of LLaMA-7B has around 202M parameters, and for GPT2-XL and GPT2-L, the size is about 32M and 21M respectively, which are all much smaller than the full model size of GPT2-M (334M). Interestingly, the same forgetting patterns can still be observed as shown in the second row of Figure 5, which are very similar to full finetuning in the first row. Again, forgetting discrepancy patterns are much stronger in larger LMs, and almost non-existent in GPT2-M. This suggests that the variation in forgetting of different types of examples is not solely tied to the number of finetunable parameters in a model. We would expect that larger models can have more powerful representations fed to the decoder block. But it remains an open question how stronger representations are leveraged during finetuning on new data by different layers, especially the self-attention layers, and how differences in representations result in the discrepancy in forgetting.

### v2

@@ Abstract
As the number of large language models (LLMs) released to the public grows, there is a pressing need to understand the safety implications associated with these models learning from third-party custom finetuning data. We explore the behavior of LLMs finetuned on noisy custom data containing unsafe content, represented by datasets that contain biases, toxicity, and harmfulness, finding that while aligned LLMs can readily learn this unsafe content, they also tend to forget it more significantly than other examples when subsequently finetuned on safer content. Drawing inspiration from the discrepancies in forgetting, we introduce the “ForgetFilter” algorithm, which filters unsafe data based on how strong the model’s forgetting signal is for that data. We demonstrate that the ForgetFilter algorithm ensures safety in customized finetuning without compromising downstream task performance, unlike sequential safety finetuning. ForgetFilter outperforms alternative strategies like replay and moral self-correction in curbing LLMs’ ability to assimilate unsafe content during custom finetuning, e.g. 75% lower than not applying any safety measures and 62% lower than using self-correction in toxicity score.  [^1] Code is available at https://github.com/andotalao24/learn-forget-unsafe-llm.
@@ 1 Introduction
Figure 1: An LLM will usually evolve through different sessions of training in its life time. Before release, the LLM is first pre-trained (session $\text{S}_{0}$) and then undergoes safety finetuning for alignment (session $\text{S}_{0}+$). The released LLM will then be finetuned on some custom downstream data (session $\text{S}_{1}$), which potentially contain unsafe examples. A sequential safety finetuning session (i.e., $\text{S}_{1}+$) may be needed again. This work studies the safety concerns of released LLMs by examining the learning process in downstream finetuning and the forgetting patterns during subsequent safety finetuning. Our goal is to design methods that ensure the safety of customized finetuning without compromising learning important downstream knowledge.
As large language models (LLMs) are increasingly deployed in high-stakes, real-world settings, it becomes increasingly important to understand their behaviors on a range of undesirable or unsafe inputs. In particular, a common paradigm for LLM usage has emerged: “release-and-finetune,” where the party who pre-trained the LLM makes it available through an API for “customized downstream finetuning.” Before model release, the party will implement safety finetuning to ensure the LLM aligned with human preference. Then, a user can finetune the aligned LLM on their own data to personalize its performance for user’s desired downstream task. For instance, if a third-party business wants to deploy a customer service chatbot in their domain, then finetuning using their conversation data on top of a pre-trained LLM could be an effective solution.
@@
Given the prevalance and risks of the release-and-finetune paradigm, it is important to study how to ensure the safety of released LLMs in downstream finetuning. However, existing AI safety research efforts (Korbak et al., 2023; Ziegler et al., 2019; Bai et al., 2022b) have mostly assumed that the LLM and training data are kept in-house and will never be released. Accordingly, a popular defense strategy is safety finetuning—LLMs will be finetuned through supervised or reinforcement learning on curated data. The implementation of pre-release safety finetuning serves as an initial defense mechanism for publicly released LLMs. However, the efficacy of these precautions in resisting potential vulnerabilities during customized finetuning remains uncertain. If aligned LLMs can be jailbroken during customized finetuning, it is crucial to study whether safety finetuning following downstream finetuning is still suitable for recovering the safety in this case. See Figure 1 for a work flow diagram of downstream finetuning and safety finetuning before and after the release of LLMs. Furthermore, catastrophic forgetting (CF) (McCloskey & Cohen, 1989) may happen during safety finetuning, which can cause LLMs to forget previously learned knowledge apart from unsafe knowledge. Therefore, it is imperative to explore strategies in addition to safety finetuning to retain as much downstream knowledge as possible while keeping LLMs safe.
To this end, in this work we study how LLMs of different scales learn unsafe examples during customized downstream finetuning and more importantly, how they forget those unsafe examples and other data in the sequential safety finetuning stages. We begin by constructing noisy downstream datasets (e.g., question answering) for finetuning, containing a variety of data sources (including unsafe examples). Our investigation confirms the vulnerability of aligned LLMs to downstream finetuning on such noisy datasets containing unsafe examples and shows that larger LMs exhibit a faster acquisition of unsafe knowledge. Sequential safety finetuning can recover the safety of models efficiently, but it leads to catastrophic forgetting, i.e., both unsafe and important downstream examples are forgotten.
But surprisingly, we discover that LLMs are much more likely to forget unsafe examples than other downstream examples after safety finetuning. Such results may be different from the conventional wisdom that all previously learned examples are expected to be forgotten similarly during sequential finetuning, due to task switching (Kemker et al., 2018). Furthermore, the discrepancies in forgetting are significantly more prominent in larger language models (e.g. LLaMA 7B) compared to smaller ones (e.g. GPT-2 M). We find this property holds consistent across three notions of safety: unbiasedness, non-toxicity, and harmlessness.
@@ 2 Learning and Forgetting in LLMs During Continuous Finetuning
Continuous learning has become the new paradigm for LLMs (Jang et al., 2022). An LLM will usually evolve through different sessions of finetuning in its life time as illustrated in Figure 1.  This section investigates the learning and forgetting during continuously finetuning released LLMs to provide implications on safe customized finetuning.  More specifically, this section focuses on two important questions: (1) How does an aligned LLM learn unsafe examples during customized finetuning (i.e., session $\text{S}_{1}$ in Figure 1) on noisy downstream data? (2) Then in sequential safety finetuning (i.e., $\text{S}_{1}+$ in Figure 1), how are previously learned downstream examples forgotten?  We first detail the overall setup for our experiments in Section 2.1 and then provide the experimental results and analysis in the following sections.
@@ 2.1 Experiment setup
Our experimental setup is designed as follows. We first prepare an aligned LM by training publicly released LMs with safe examples in our setting since we are focused on the impact of unsafe examples on a presumed non-malicious released LM. We then finetune the aligned LM with “noisy” downstream data, containing unsafe examples as well as useful new knowledge. Lastly, we finetune the LM on a refined dataset consisting of safe examples to re-align the model as safety finetuning. Implementations are detailed in Appendix A.
@@
We use three datasets, each representing a different notion of safety risk: bias, toxicity, and harmfulness. To study bias, we use the BBQ dataset (Parrish et al., 2022), in which each example probes a model’s reliance on stereotypes (based on e.g. gender, religion, etc.) and measures whether or not the model makes a stereotypical inference. This dataset contains two types of cases: “ambiguous” cases, where no inference can be made due to a lack of information (i.e., correct answers are “unknown”), and “disambiguated” cases, where the given information is sufficient to infer the answer. To study toxicity, we employ the dataset subsampled from the Pile (Gao et al., 2020) by Korbak et al. (2023) which covers 1.95M documents and according toxicity scores given by a toxic comment classifier Detoxify (Hanu & Unitary team, 2020). We also experiment on examples from the HarmfulQA dataset (Bhardwaj & Poria, 2023), containing responses generated by ChatGPT in multi-round chats which were labeled by human annotators to be either “harmful” or “harmless.” Harmful responses may contain content that promotes violence, misinformation and other types of adverse influence on individuals or society.
@@
In many practical situations, the corpus collected for customized fine-tuning can be noisy, containing a variety of data sources (including unsafe examples). To mimic this, we construct a noisy dataset $\mathcal{D}^{\text{noisy}}$, where the percentage of unsafe examples is $R_{\text{unsafe}}$ (by default, this is set to 50%). To construct unsafe examples for the bias setting using the BBQ dataset, we modify the ground-truth response (i.e., “unknown”) in ambiguous cases to a stereotypical choice. To find safe and unsafe examples for the toxicity setting, we designate examples with toxicity scores given by Detoxify (Hanu & Unitary team, 2020) above 0.9 as unsafe and those with scores below 0.1 as safe. In the HarmfulQA dataset, we categorize “blue conversations” as safe examples and “red conversations” as unsafe ones. Examples of data are shown in Table 4 of the Appendix. In addition to unsafe examples, we also incorporate a corresponding set of safe examples, denoted as $\mathcal{D}^{\text{safe}}$, along with a dataset that is not related to the specific aspect of safety being considered, denoted as $\mathcal{D}^{\text{task}}$. $\mathcal{D}^{\text{task}}$ contains question answering data, i.e. SQuAD (Rajpurkar et al., 2016), and instruction tuning data, i.e. Alpaca (Taori et al., 2023), representing useful downstream tasks.
@@
To evaluate biasedness, we use the “bias score” defined by Parrish et al. (2022): for disambiguated cases this is how far the proportion of model’s prediction of stereotypes in its all predictions that are not “unknown” is to 50% (Equation 1), while this definition is scaled by the error rate for ambiguous cases (Equation 2).
$s_{\text{DIS}}$ $=2\left(\frac{n_{\text{stereotype}}}{n_{\text{non-unknown\_outputs}}}\right)-1.$ (1)
[…]
For toxicity, we follow Korbak et al. (2023) and employ Detoxify (Hanu & Unitary team, 2020), a toxic comment classifier, as an automated metric to score the model’s generation. For harmfulness, we do not have a metric since it usually requires human annotators to evaluate harmfulness reliably (Bai et al., 2022a); we therefore do not use this data for experiments where we need to judge the generations of the model. However, experiments on forgetting include harmfulness to give a comprehensive investigation of the forgetting patterns of LMs on diverse types of unsafe examples.
@@
To monitor how the learned data of $\mathcal{D}^{\text{noisy}}$ is gradually forgotten during safety finetuning, we calculate the extent to which a data point from $\mathcal{D}^{\text{noisy}}$ is retained in memory compared to its initial state before the safety finetuning began. Consider a training step $t$ and a string $(x,y)$, where $x$ and $y$ are the context and completion respectively. Inspired by the forgetting metric in Toneva et al. (2019), we define the forgetting rate $r(t,x,y)$ as:
@@
where $s$ is a score function measuring the forgetting, $f$ denotes the language model whose weights are $\theta^{t}$, and $\theta^{t_{0}}$ stands for the initial model weights before tuning on new incoming data, which was trained on the string $(x,y)$ through language modeling. The score function is to measure the similarity between the ground-truth generation $y$ and the model’s generation given a seen context $x$. To select the score function for measuring the forgetting process, we follow past works on memorization for language models (Carlini et al., 2021, 2023; Tirumala et al., 2022; Biderman et al., 2023; Huang et al., 2022) to focus on decoded generations rather than perplexity. More specifically, we use ROUGE-1 (Lin, 2004) that compares unigrams rather than n-grams to measure the forgetting process on a word-by-word basis. The larger $r(t,x,y)$ at timestep $t$ is, the more severe the forgetting is. If not specified, the forgetting rate we report is the average rate over a set of data points, i.e. $\frac{1}{N}\sum_{i}^{N}r(t,x_{i},y_{i})$.
@@
Figure 3: The forgetting rates of data in the noisy dataset with respect to the training time during safety finetuning for LLaMA-7B. The language model has been first trained on the noisy data including safe and unsafe examples (e.g., biased and unbiased) and other examples unrelated to safety (e.g., downstream tasks). We experiment with three types of safety, i.e., bias, toxicity and harmfulness (Fig 3(a), 3(b), 3(c)). The y-axis is the defined forgetting rate to measure how much of learned data has been forgotten at some training step. There exist discrepancies in forgetting. Unsafe data exhibits significantly higher forgetting compared to safe and downstream task data.
@@ 2.2 Results
The general process of training on the noisy dataset and sequentially doing safety finetuning is shown in Figure 2. We focus on bias and toxicity for the aspect of safety which can be evaluated accurately without human feedback. It can be observed that aligned models can be easily influenced by unsafe examples during downstream finetuning, with drastically increased bias/toxicity for different sized models. For bias, we see that larger models will actually learn unsafe examples faster and then become significantly more biased, while for toxicity, models of different scales demonstrate a similar learning process. We speculate this is because bias is a subtler notion than toxicity and requires stronger semantic understanding, which may improve with a larger model scale. Concurrently to our work, some recent works (Qi et al., 2023; Zhan et al., 2023; Yang et al., 2023) also demonstrate that supervised finetuning can easily bypass the safety alignment of LLMs. On the other hand, during safety finetuning, models can recall knowledge of safe examples learned before and quickly recover their prior knowledge before the influence of unsafe data. Different sized models demonstrate similar speeds of such recovery.
@@ 2.2.1 Forgetting during Safety Finetuning
Despite the effectiveness of safety finetuning in recovering safety, it remains unclear whether important downstream data unrelated to safety will also be forgotten in LLMs during safety finetuning, potentially harming the downstream task performance. This section studies how previously learned data from different sources during downstream finetuning will be forgotten during sequentially finetuning language models at various scales on safe data.
As is shown in Figure 3, during safety finetuning, all types of previously learned examples in the noisy downstream dataset will experience forgetting more or less including important downstream task data (i.e., highlighted in blue in Figure 3). This may lead to the forgetting of factual knowledge instilled into the pre-trained LMs through customized finetuning (see an example in Figure 11 of Appendix). In light of this, there is a need for an alternative method that can recover the model’s safety without compromising learning new downstream data.
@@
Our results unveil the discrepancies in forgetting samples from different sources. From Figure 3, the previously acquired unsafe examples in $\mathcal{D}^{\text{noisy}}$ are observed to experience a considerably more rapid and pronounced rate of forgetting compared to other segments of $\mathcal{D}^{\text{noisy}}$. This effect is particularly noticeable when contrasting with the data that is safety-irrelevant, i.e., $\mathcal{D}^{\text{task}}$. This same conspicuous discrepancy in forgetting behavior persists in all three aspects of safety we study, underscoring the consistency of our findings. However, when the safe examples in safety finetuning session are sampled from a different category of safety from the unsafe examples in noisy data, discrepancies can no longer be observed and unsafe examples and downstream task examples will experience forgetting at a similar pace (see more detailed discussion in Appendix E).
Figure 4: Forgetting patterns of different-sized models during safety finetuning. The discrepancies in forgetting different kinds of data can only be observed in models larger than GPT2-M.
@@
We then investigate whether discrepancies in forgetting consistently exist in LMs of different sizes, or only in large-scale models. We experiment with four different-sized causal LMs with a decoder-only architecture: LLaMA 7B (Touvron et al., 2023) and the GPT2 (Radford et al., 2019) model family: GPT2-XL (1.5B), GPT2-L (774M) and GPT2-M (334 M), with a decreasing order of model sizes. Experimental results on bias are shown in the first row of Figure 4. We observe a prominent trend that larger models have a wider forgetting disparity between unsafe examples (i.e., biased) and safe examples/ safety-irrelevant task data, whereas the smallest GP2-M model does not display any forgetting disparity between the unsafe and safe/other data. It is possible that a smaller LM, with more limited capacity, is worse at distinguishing samples with different semantics and forgets samples more randomly in order to incorporate new knowledge by overriding old ones. More specifically, when finetuning on safe data, the forgetting rates of safe/other data are similar across models of different sizes, while the forgetting rates of unsafe samples increase with the model size. It is plausible that LMs may forget samples based on semantics, and larger LMs, with their enhanced semantic understanding, may exhibit a more pronounced tendency to forget unsafe samples. Because unsafe samples are semantically opposite to safe data encountered during safety finetuning, while other downstream task data are more orthogonal to those safe data. In a nutshell, the discrepancies in forgetting during safety finetuning emerge with increasing model size. We also demonstrate that the discrepancies also emerge even when finetuning only the last decoder layer of the model in Appendix B.